OpenStudy (anonymous):

–6q + 5.8 = 23.8 A) q = 3 A) q = –3 A) q = –1.83 A) q = 4.93 Can you teach my how to do it i'm not to sure about this one Thanks for help!

OpenStudy (anonymous):

–6q + 5.8 = 23.8 1st step, do you know how to move the 5.8 to the other side?

OpenStudy (anonymous):

do we divide 5.8 by both sides? first

OpenStudy (anonymous):

I'm going by my other problem but it doesnt seem to look like this one

OpenStudy (anonymous):

no you have to subtract 5.8 from both sides

OpenStudy (anonymous):

oh ok so –6q + 5.8 = 23.8 -5.8 -5.8 -6q + 0 = 18

OpenStudy (anonymous):

then do I minus 6 by each side?

OpenStudy (anonymous):

now you have -6q = 18 now you need to divide

OpenStudy (anonymous):

-1q = 3

OpenStudy (anonymous):

you need to divide by the -6 so you get q = -3